Lane Company manufactures a single product and applies overhead cost to that product using standard direct labor-hours. The budgeted variable manufacturing overhead is $2.80 per direct labor-hour and the budgeted fixed manufacturing overhead is $612,000 per year.

The standard quantity of materials is 4 pounds per unit and the standard cost is $5.00 per pound. The standard direct labor-hours per unit is 1.5 hours and the standard labor rate is $12.40 per hour.

The company planned to operate at a denominator activity level of 90,000 direct labor-hours and to produce 60,000 units of product during the most recent year. Actual activity and costs for the year were as follows:

Actual number of units produced 72,000
Actual direct labor-hours worked 117,000
Actual variable manufacturing overhead cost incurred $ 210,600
Actual fixed manufacturing overhead cost incurred $ 643,500

Required:

1. Compute the predetermined overhead rate for the year. Break the rate down into variable and fixed elements.

2. Prepare a standard cost card for the company’s product.

3a. Compute the standard direct labor-hours allowed for the year’s production.

3b. Complete the following Manufacturing Overhead T-account for the year.

4. Determine the reason for any underapplied or overapplied overhead for the year by computing the variable overhead rate and efficiency variances and the fixed overhead budget and volume variances.
Answer #1

1.

Computation of Predetermined overhead rate - Lane Company
Particulars Amount
Budgeted variable manufacturing overhead ($2.80*90000) $2,52,000
Budgeted fixed manufacturing overhead $6,12,000
Total Manufacturing overhead $8,64,000
/Budgeted direct labor hours 90000
Predetermined overhead rate $9.60
Less: Variable overhead rate (Per direct labor hour) $2.80
Fixed manufacturing overhead rate (per direct labor hour) $6.80

2.

Standard Cost Card - Lane Company
Direct Material 4 Pounds at $5.00 per pound $20.00
Direct labor 1.5 DLHs at $12.40 per DLH $18.60
Variable manufacturing overhead 1.5 DLHs at $2.80 per DLH $4.20
Fixed manufacturing overhead 1.5 DLHs at $6.80 per DLH $10.20
Standard cost per unit $53.00

3a.

Standard hours allowed = 72000 *1.5 = 108,000 hours

3.

Manufacturing Overhead
Particulars Debit Particulars Credit
To Cash (Variable manufacturing overhead) $2,10,600 By WIP (Applied overhead) (108000*$9.60) $10,36,800
To Cash (Fixed manufacturing overhead) $6,43,500
To overhead variance (Overapplied overhead) $1,82,700
Total $10,36,800 Total $10,36,800

4.

Variable overhead actual rate = $210600/ 117000 = $1.80 per hour

Variable overhead rate variance = (SP - AP) *Actual Hours = ($2.80 - $1.80) * 117000 = $117,000 Favorable

Variable overhead Efficiency variance = (Standard hours - Actual Hours) *SP = (108000 - 117000)*$2.80 = - $25,200 (Unfavorable)

Fixed overhead budget Variance = Budgeted Fixed Overhead - Actual Fixed overhead = $612000 - $643500 = -$31,500 (unfavorable)

Fixed Overhead Volume Variance = ($6.80*108000) - $612000 = $122,400 Favorable